An accounting firm has recently recruited six graduates: three men and three women. Three of the graduates are to be selected at random to work in the firm’s suburban office. a. Construct the probability tree. b. What is the probability that at least two women will be selected? c. If we know that at least two women were selected what is the probability that all the selected graduates are women?
Answer #1

(a) Here the probability tree

(b) Pr(at least two women would be selected ) = Pr(there are 2 women out of 3 women) + Pr(there are 3 women out of 3 women)

= 3C2* 3C1/6C3 + 3C33C0/6C3

= 0.45 + 0.05 = 0.50

(c) Pr(all the selected graduates are women given at least two women were selected) = 0.05/(0.45 + 0.05) = 0.10